# Machine-generated text detection prevents language model collapse

### v1

@@ 6.2 Results
Our method is evaluated against the baseline of training on all samples in the pool of data, $\mathcal{D}$. In addition, we define the “Oracle” performance which represents a perfect AI text detector that filters all synthetic samples. The task performance metrics, data quality metrics and detector accuracy over generations of partially synthetic training $\left(\alpha=1,\beta=1,\gamma=0\right)$ can be seen in Figure 2. We also display the performance in the setting with half the human data $\alpha=0.5$ in Appendix S1. Over $9$ iterations of recursive training, our method of weighted sampling can be seen to prevent model collapse and when $\alpha=1$ i.e. there is an equal proportion of human to synthetic data, our method improves model performance compared to training on exclusively human data (Oracle) across both models and respective decoding strategies. In addition to the improved task performance, our method can also be seen to remove any effects on the readability, diversity and detectability of the synthetic samples over generations compared to the baseline. Improving performance on the Oracle demonstrates the value of synthetic data in LLM training. However, without our method resampling the distribution to contain more human-generated samples and removing likely synthetic text, the model collapses.
@@
Figure 3: Perplexity Distribution for machine-generated text of GPT-2 at generations 0, 1 and 9 in the partially synthetic scenario ($\alpha=0.5$). The graphs on the left are the baseline and the graphs on the right are when our method is used. Perplexity is evaluated using the model trained on the human text. Machine-generated text from GPT-2 using top-$k$ decoding and sampling.
We evaluate recursive training under three different mixing coefficient settings. In Table 3 we present the percentage difference between the final model under the baseline strategy and using our method. From the table, it can be seen that mixing cross-generational data actually has an insignificant effect on the extent of model collapse compared to training with previous generational data only, contrasting with the findings of Gerstgrasser et al. (2024). However, in their experiments, the compute budget was unlimited whereas we take the more realistic setting of creating a fixed-size dataset using mixing coefficients.
@@
We evaluate the perplexity distribution of the synthetic data at each generation using the model $p^{0}$ trained on the human data, similar to Shumailov et al. (2024). We plot in Figure 3 the distributions for generation 00, $1$ and $9$ for the GPT-2 model with top-$k$ and pure sampling and compare our method to the baseline. For top-$k$ decoding, we observe a similar behaviour to the results of Shumailov et al. (2024) in which the perplexity shifts towards regions of lower perplexity and the distribution becomes more peaked, showing signs of early model collapse. For pure sampling, on the other hand, we see a completely different behaviour, the distributions shift right and become more spread out into regions of low probability and high perplexity. This is an interesting finding that demonstrates that by removing truncation from the decoding strategy, the narrowing effect of model collapse is removed and instead, we see the model collapse in a completely different way by generating long-tail incoherent text that is completely distinct from the original human samples. Using our mitigation strategy, however, we observe very little change in the perplexity distribution for both sampling strategies and improved data quality metrics to the baseline, as seen in Table 3.
@@ 7 Conclusion
In this work we investigate model collapse across three dimensions: model performance, data quality and resemblance to human samples. Through our analysis we found that the extent of model collapse and the effect on the data distribution is influenced by the decoding strategy where truncating the distribution can lead to peaked distributions and repetitive models while sampling directly can result in high perplexity and verbose outputs with low resemblance to human data. Using the decoding strategies that resulted in the most extreme collapse, we evaluated the partially synthetic scenario where human data is mixed into the training data, but in an unknown proportion. We designed a novel method to mitigate model collapse based on resampling the training distribution using the predictions of a machine-generated text detector. We have validated our method on two popular LLMs (GPT-2 and SmolLM2) across a range of decoding strategies, showing that we can prevent model collapse in all cases. When the proportion of human data is high enough, our method results in improved model performance compared to training only on the human data.
@@
The primary limitation of this study is the evaluation of relatively small LLMs (137M and 360M parameters) compared to larger models, which often have billions of parameters, limiting the generalisability of these findings. However, recent work Dohmatob et al. (2024a) suggests that model collapse is even more pronounced in larger models, indicating that our proposed approach could be even more beneficial in such cases. Future work could extend this analysis to larger models to validate these findings.
Additionally, similar to previous studies Shumailov et al. (2023); Dohmatob et al. (2024b), we assess LLMs exclusively in a fine-tuning setting rather than pretraining from scratch. While pretraining experiments could provide deeper insights, the computational cost and complexity of training large-scale models from the ground up make such an approach impractical in our case. Nevertheless, given that model collapse has been primarily evaluated in LLMs from a finetune setting, the conclusions made in this work still align with the current body of research.
@@
Appendix A AI text detection
A.1 Pretrained models
RoBERTa
(Robustly Optimized BERT Pretraining Approach) Goyal et al. (2020) improves on the pre-training phase of BERT Kenton and Toutanova (2019), an encoder-only Transformer model that leverages masked language models to enable pre-trained deep bidirectional representations. The RoBERTa model optimised the pre-training procedure for BERT by training the model for longer and on more data, changing the masking pattern, and removing the next sentence prediction objective. We use the RoBERTa base variant which has $125$ million parameters.
DeBERTav3
(Decoding-enhanced BERT with Disentangled Attention) He et al. (2023), is a BERT-based encoder only model enhanced with disentangled attention. DeBERTav3 model improves on DeBERTa by using Enhanced Mask Decoding and an ELECTRA-style pretraining objective, Replaced Token Detection, instead of Masked Language Modeling. We use the base variant which contains $86$ million backbone parameters with an embedding layer of $98$ million parameters.
ModernBERT
Warner et al. (2024), is a recent addition to the encoder-only transformer models that has been designed to increase downstream performance and efficiency on GPUs, particularly in long-context scenarios due to its 8192 native sequence length. The model was trained on 2 trillion tokens and improves on the original BERT architecture with rotary positional embeddings (RoPE), unpadding, GeGLU layers and alternating local-global attention demonstrating state-of-the-art performance amongst encoder models across a range of classification and retrieval tasks. We conduct experiments with the base variant, which contains $150$ million parameters.
A.2 Dataset
@@
The data used for the human text consists of opinion statements (CMV & Yelp reviews dataset), news articles (XSum & TLDR dataset), question answering (ELI5), story generation (Reddit WritingPrompts & ROC), commonsense reasoning (HellaSwag), knowledge illustration (SQuAD) and Scientific writing (SciGen). They sample 1000 texts from each domain (apart from opinion statements and news articles with 804 and 777 samples respectively) and generate text from 27 LLMs from 7 model families, which include OpenAI, LLaMA, GLM, FLAN-T5, OPT, BigScience and EleutherAI. For each human-written sample in the dataset, they generate a machine-generated version by providing the LLM the first 30 tokens of human-written text. In addition, for the OpenAI models, they investigate two other prompt strategies for applicable domains. These include ’topical’ prompts such as an argument or news title and ’specified’ prompts which contain information about the domain source. This results in 33,000 (27,000+3*2*1000) machine-generated samples per source before processing and filtering. The authors split the dataset into train, validation and test splits in the ratio $80:10:10$. To mitigate data imbalance in the validation and test sets they sample additional human data from each data source so that both sets have a roughly 50:50 balance. The training set, however, is $71\%$ machine-generated. All splits combined amount to $154,078$ human-written and $294,381$ machine-generated texts. In addition to the test set, we also evaluate our detector on their challenging test set containing text from four unseen domains (CNN/DailyMail, DialogSum, PubMedQA, IMDB) and generated by an unseen model (GPT4).
@@
GPT-2 Greedy 29.29/82.74 38.72/34.93 0.96/0.70 61.02/67.13 0.99/1.00 60.47/8.25
Beam Search 29.25/78.06 38.75/35.21 16.78/10.86 61.54/67.60 0.91/1.29 60.97/17.57
Sampling 29.29/58.64 38.74/32.49 94.88/99.82 24.12/6.76 90.16/7.18 40.62/-10.14
Temperature 29.23/44.55 38.77/34.47 87.76/25.10 33.45/54.56 94.15/22.69 46.80/36.79
Top-$k$ 29.31/48.36 38.73/33.12 84.57/70.20 38.81/42.14 95.21/70.01 51.19/34.32
Nucleus 29.28/48.96 38.74/32.73 92.26/86.73 28.24/26.86 90.96/57.41 43.96/21.31
SmolLM2 Greedy 13.36/90.24 47.89/44.00 6.99/3.02 61.90/50.23 4.29/1.12 62.33/46.72
Beam Search 13.36/88.90 47.87/43.96 6.75/2.88 61.47/44.47 3.86/1.09 62.38/44.47
Sampling 13.36/14.76 47.88/46.82 90.83/91.62 37.69/33.33 91.28/83.82 47.49/44.33
Temperature 13.36/23.70 47.86/46.38 82.96/24.37 46.21/61.64 91.48/21.66 52.78/64.59
Top-$k$ 13.36/19.01 47.88/46.41 82.57/57.35 48.62/61.18 95.48/68.41 55.45/62.82
Nucleus 13.36/21.80 47.88/46.55 87.38/47.36 42.43/60.85 94.40/67.37 51.14/63.71
Human 88.79 42.89 100 50.34
Table S1: Performance results and qualitative scores for GPT-2 and SmolLM2 over a range of decoding strategies. Generation 0 and 9 are presented and recursive training is in the fully synthetic scenario.
A.3 Hyperparameters
Each model was finetuned for 5 epochs with an early stopping patience of 3 epochs which is based on the macro average F1 score on the validation set and with no warm up. Models were evaluated on the validation set after every epoch which was used to select the best model for evaluation. Optimization was performed using AdamW with a $\beta_{1}$, $\beta_{2}$, $\epsilon$ and weight decay of $0.9$, $0.98$, $1e^{-6}$ and $1e^{-2}$ respectively. These parameters were chosen based on the recommended parameters of Warner et al. (2024). The label smoothing parameter $\alpha$ was set to 0.1. The seed was fixed at $42$ and a training batch size of $16$ and validation and test batch size of $64$ was used for all experiments. The learning rate was set based on a hyperparameter sweep over [$1e^{-5}$, $1.5e^{-5}$, $2e^{-5}$, $3e^{-5}$, $4e^{-5}$]. For ModernBERT the best performing learning rate was $1e^{-5}$. The classification threshold was selected to maximise the macro average F1 score on the validation set.
Appendix B Model collapse experiments
B.1 Model collapse training
In our experiments, we conduct full fine-tuning on the GPT-2 base (128 million parameters) and SmolLM2 (360 million parameters) using a learning rate of $5\!\times\!10^{-5}$, batch size of $8$ and a dropout rate of $0.1$. For the AdamW optimizer, we use $\beta_{1}$, $\beta_{2}$ and $\epsilon$ of $0.9$, $0.999$ and $081e-$ 08 respectively. Each model is trained for $1$ epoch with the hyperparameters fixed for all experiments. We conduct 10 iterations of recursive training.
B.2 Open-ended text generation dataset
All our model collapse experiments use the raw variant of the WikiText-2 dataset Merity et al. (2016). We train models on the training set, consisting of $36718$ documents and evaluate on the test set of $4358$ documents. The WikiText dataset was extracted from the Good or Featured article criteria specified by editors on Wikipedia and only covers the English language.
B.3 Models
We conduct experiments on the base variant of GPT-2 ($137$ million parameters) and the $360$ million parameter variant of SmolLM2.
Figure S1: GPT2 (top) and SmolLM2 (bottom) under partially synthetic recursive training (alpha=0.5) for 10 generations. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect AI text detector that filters all synthetic samples.
[…]
GPT-2 Top-$k$ 0.5 1 0 84.46 / 82.16 39.10 / 39.55 93.31 / 91.10 51.30 / 46.16
0.5 0.5 0.5 85.66/84.35 38.94/40.64 94.07/92.00 51.06/45.38
1 1 0 84.66 / 82.68 39.19 / 40.20 93.24 / 92.69 50.99 / 48.09
Sampling 0.5 1 0 94.86 / 97.77 24.09 / 14.52 91.08 / 46.87 40.64 / 23.69
0.5 0.5 0.5 94.95/97.50 24.15/17.02 90.07/76.01 41.02/29.08
1 1 0 94.88 / 96.86 24.04 / 18.12 91.18 / 75.64 40.71 / 31.49
GPT-2 mitigation Top-$k$ 0.5 1 0 84.43 / 85.89 39.43 / 38.85 94.59 / 91.52 50.79 / 50.33
0.5 0.5 0.5 85.18/86.40 39.28/38.90 94.75/95.67 51.31/49.99
1 1 0 84.81 / 85.65 38.85 / 38.76 94.92 / 93.95 51.28 / 51.34
Sampling 0.5 1 0 94.87 / 96.50 24.30 / 19.85 92.92 / 81.58 40.68 / 35.59
0.5 0.5 0.5 94.92/96.46 24.09/20.66 91.16/88.46 40.98/36.48
1 1 0 94.61 / 96.17 24.01 / 21.87 91.76 / 88.06 40.78 / 38.10
SmolLM2 Top-$k$ 0.5 1 0 82.50/73.99 48.99/52.77 94.58/91.10 55.65/58.09
0.5 0.5 0.5 82.46/76.24 48.66/52.44 93.68/95.13 55.24/57.88
1 1 0 82.86/77.80 48.68/51.36 94.76/89.06 55.43/57.27
Temperature 0.5 1 0 83.42/66.25 45.57/52.77 95.37/81.27 52.72/57.67
0.5 0.5 0.5 83.27/71.22 45.89/52.73 88.33/84.54 53.01/56.23
1 1 0 82.73/72.75 45.69/52.21 94.57/87.55 52.88/56.39
SmolLM2mitigation Top-$k$ 0.5 1 0 82.21/81.77 49.04/49.91 94.96/93.19 55.49/56.64
0.5 0.5 0.5 82.56/81.86 48.83/49.62 94.55/94.46 55.33/56.55
1 1 0 82.35/82.52 45.00/48.91 94.00/94.43 55.35/56.14
Temperature 0.5 1 0 82.83/79.35 46.08/49.15 93.13/70.63 52.87/56.08
0.5 0.5 0.5 83.31/79.62 45.64/49.34 93.07/96.14 52.79/55.68
1 1 0 83.24/80.87 45.94/48.50 93.93/78.36 52.64/54.90
@@ B.3 Models
Table S2: Comparison of data quality when using our mitigation strategy versus the baseline. Results are shown for top-k decoding and pure sampling with different values for $\alpha$, $\beta$, and $\gamma$. The proportion of human data in the recursive training is given by $\alpha$. $\uparrow$ means higher is better and $\downarrow$ means lower is better.

### v2

@@ 5.3 Results on collapse prevention
We evaluate recursive training in the partially synthetic setting under $3$ mixing coefficient settings ($\alpha\!=\!1$, $\beta\!=\!1$, $\gamma\!=\!0$, $\alpha\!=\!0.5$, $\beta\!=\!1$, $\gamma\!=\!0$ and $\alpha\!=\!\beta\!=\!\gamma\!=\!0.5$). The task performance, data quality, and detector accuracy metrics over $9$ generations of partially synthetic training are depicted in Figure 2 for the scenario when there is an equal proportion of human and synthetic samples (we also provide results for the $2$ additional scenarios in Appendix S2,S3 ). Weighted sampling (as described in section 5.2) appears to prevent model collapse, while the baseline degrades (higher perplexity, lower test accuracy). Our method also alleviates any effects on the readability and diversity of the synthetic samples. The detectability of the baseline text, on the other hand, is increasing as it becomes more dissimilar to human text. Interestingly, our method improves performance compared to training exclusively on human data (Oracle) for both language models and respective decoding strategies. These outcomes demonstrate both the value of using synthetic data in LLM training, but also the importance of selecting the right synthetic samples while maintaining a high proportion of human data.
@@
Figure 3: Distribution of perplexity for machine-generated text of GPT-2 at generations 00, $1$, and $9$ in the partially synthetic scenario $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$. Perplexity is evaluated using the model trained on the human text ($p^{0}$).
Table 3 enumerates the percentage difference for the final models under the baseline strategy vs. using our approach. We also enumerate the exact scores in Table S1. Our method improves the model output quality across all metrics compared to the baseline except for sampling with GPT-2. In this case, the diversity of the text increases dramatically using the baseline method, however, this is accompanied by a significant drop in MAUVE and readability compared to our method. We observe that mixing cross-generational data has an insignificant effect on the extent of model collapse compared to training solely on the previous generation, contrasting with the findings of Gerstgrasser et al. (2024). However, we note that Gerstgrasser et al. (2024) did not constrain the sample size and trained on all data available. Our experiments adopt the more realistic and less extreme setting, sampling the pool of data under mixing scenarios.
@@
Similar to Shumailov et al. (2024), we also study the perplexity distribution of the synthetic data at each generation using model $p^{0}$ that was trained on human data. Figure 3 depicts these distributions for generations 00, $1$, and $9$ for the GPT-2 model with top-$k$ and pure sampling, compared to the baseline. For top-$k$ decoding, similarly to Shumailov et al. (2024), we observe that for the baseline perplexity shifts towards regions of lower scores and the distribution becomes more peaked, showing signs of early model collapse. For pure sampling, on the other hand, we observe that the baseline distribution shifts to higher perplexity scores and becomes more spread out. This is an interesting finding that demonstrates that by removing truncation from the decoding strategy, the narrowing effect of model collapse is diminished and instead, model collapse is reflected by long-tail incoherent text that is completely distinct from the original human samples. By deploying our mitigation strategy, however, we observe very little change in the perplexity distribution for both sampling strategies, and also improved data quality metrics to the baseline, as seen in Table 3.
@@ 6 Conclusion
This work investigates model collapse across three dimensions: model performance, data quality, and resemblance to human samples. Through our analysis we found that the extent of model collapse and the effect on the data distribution is influenced by the decoding strategy. Truncating can lead to peaked distributions and repetitive models while sampling directly can result in high perplexity and verbose outputs with low resemblance to human data. Using the decoding strategies that resulted in the most extreme collapse, we evaluated the partially synthetic scenario, where human data is mixed into the training data. We designed a novel method to mitigate model collapse based on resampling the training distribution using the predictions of a machine-generated text detector. We have validated our method on two popular LLMs (GPT-2 and SmolLM2) across a range of decoding strategies, showing that we can prevent model collapse in all cases. When there is an equal ratio of human to synthetic samples in the training pool, our method results in improved model performance compared to training only on the human data.
@@
The primary limitation of this study is the evaluation of relatively small LLMs ($117$M and $360$M parameters) compared to larger models, which often have billions of parameters, limiting the generalisability of these findings. However, recent work Dohmatob et al. (2024a) suggests that model collapse is even more pronounced in larger models, indicating that our proposed approach could be even more beneficial in such cases. Future work could extend this analysis to larger models to validate these findings.
Additionally, as in previous studies Shumailov et al. (2023); Dohmatob et al. (2024b), we assess LLMs exclusively in a fine-tuning setting rather than pre-training from scratch. While pre-training experiments could provide deeper insights, the computational cost and complexity of training large-scale models from the ground up make such an approach impractical in our case. Nevertheless, given that model collapse has been primarily evaluated in LLMs from a fine-tuning setting, the conclusions made in this work still align with the current body of research.
@@
Acknowledgments
G. Drayson is funded by the EPSRC grant “AI Centre for Doctoral Training in Foundational Artificial Intelligence” (EP/S021566/1). V. Lampos would like to acknowledge the support from the EPSRC grant “Digital Health Hub for Antimicrobial Resistance” (EP/X031276/1).
@@ Acknowledgments
Appendix A Machine-generated text detection
A.1 Dataset
@@
We trained and evaluated the machine-generated text detectors on the MAGE dataset Li et al. (2024), which is based on documents from $10$ domains: opinion statements (CMV & Yelp reviews dataset), news articles (XSum & TLDR dataset), question answering (ELI5), story generation (Reddit WritingPrompts & ROC), commonsense reasoning (HellaSwag), knowledge illustration (SQuAD) and Scientific writing (SciGen). The authors sampled $1{,}000$ texts from each domain (apart from opinion statements and news articles with $804$ and $777$ samples respectively) and generated text using $27$ LLMs from $7$ model families, which include OpenAI, LLaMA, GLM, FLAN-T5, OPT, BigScience and EleutherAI. For each human-written sample in the dataset, they generate a machine-generated version by providing the first $30$ tokens of human-written text as context to the LLM. In addition, for the OpenAI models, they implemented two other prompt strategies for relevant domains: ‘topical’ prompts such as an argument or news title and ‘specified’ prompts which contain information about the domain source. This results in $33{,}000$ ($27{,}000\!+\!3\times 2\times 1{,}000$) machine-generated samples per source before processing and filtering. The authors split the dataset into train, validation and test splits in the ratio $80:10:10$. To mitigate data imbalance in the validation and test sets they sample additional human data from each data source. The resulting test set contains $28{,}741$ human and $28{,}078$ machine-generated samples ($49\%$ machine-generated). The training set, however, is $71\%$ machine-generated. The total dataset consists of $154{,}078$ human-written and $294{,}381$ machine-generated texts. In addition to the test set, we also evaluate our detector on their more challenging test set containing text from four unseen domains (CNN/DailyMail, DialogSum, PubMedQA, IMDB) and generated by an unseen model (GPT-4). This out-of-distribution test set contains $762$ human and $800$ machine-generated samples.
Figure S1: Wikitext test set performance (perplexity and accuracy) for varying mixing coefficients ($\alpha$, $\beta$, $\gamma$) across generations 00 to $9$ using pure sampling decoding with GPT-2.
GPT-2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SmolLM2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure S2: GPT-2 (top) and SmolLM2 (bottom) under partially synthetic recursive training $\left(\alpha\!=\!0.5,\beta\!=\!1,\gamma\!=\!0\right)$ for 10 generations. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect AI text detector that filters all synthetic samples.
A.2 Pre-trained models
Robustly Optimized BERT pre-training Approach (RoBERTa) by Liu et al. (2019) improves on the pre-training phase of BERT Devlin et al. (2019), an encoder-only transformer model that leverages masked language models to enable pre-trained deep bidirectional representations. The RoBERTa model optimised the pre-training procedure for BERT by training the model for longer and on more data, changing the masking pattern, and removing the next sentence prediction objective. We use the base variant which has $125$M parameters.
Decoding-enhanced BERT with Disentangled Attention (DeBERTav3) by He et al. (2023), is a BERT-based encoder only model enhanced with disentangled attention. DeBERTav3 model improves on DeBERTa by using Enhanced Mask Decoding and an ELECTRA-style pre-training objective, Replaced Token Detection, instead of Masked Language Modelling. We use the base variant which contains $86$M backbone parameters with an embedding layer of $98$M parameters.
ModernBERT Warner et al. (2024) is a recent addition to the encoder-only transformer models that has been designed to increase downstream performance and efficiency on GPUs, particularly in long-context scenarios due to its $8{,}192$ native sequence length. The model was trained on $2$ trillion tokens and improves on the original BERT architecture with rotary positional embeddings (RoPE), unpadding, GeGLU layers and alternating local-global attention demonstrating SOTA performance amongst encoder models across a range of classification and retrieval tasks. We conduct experiments with the base variant, which contains $150$M parameters.
A.3 Hyperparameters
Each model was fine-tuned for 5 epochs with an early stopping patience of 3 epochs which is based on the macro average F1 score on the validation set and with no warm-up. Models were evaluated on the validation set after every epoch which was used to select the best model for evaluation. Optimisation was performed using AdamW with a $\beta_{1}$, $\beta_{2}$, $\epsilon$, and weight decay of $0.9$, $0.98$, $10^{-6}$, and $10^{-2}$, respectively. These parameters were chosen based on the recommended parameters of Warner et al. (2024). The label smoothing parameter $\alpha$ was set to 0.1. The seed was fixed at $42$ and a training batch size of $16$ and validation and test batch size of $64$ were used for all experiments. The learning rate was set based on a hyperparameter sweep over [$1$, $1.5$, $2$, $3$, $4$]×$\times 10^{-5}$. For ModernBERT the best-performing learning rate was $10^{-5}$. The classification threshold was selected to maximise the macro average F1 score on the validation set.
Appendix B Model collapse experiments
B.1 Dataset
All our model collapse experiments use the raw variant of the WikiText-2 dataset Merity et al. (2016). We train models on the training set, consisting of $36,718$ documents and evaluate on the test set of $4,358$ documents. The WikiText dataset was extracted from the Good or Featured article criteria specified by editors on Wikipedia and only covers the English language.
B.2 LLMs
The Generative Pre-trained Transformer 2 (GPT-2Radford et al. (2019) is a decoder-only transformer-based language model. GPT-2 demonstrated that large-scale language models could perform various language tasks without task-specific training. We use the base variant, which contains $117$M parameters. SmolLM2 Allal et al. (2025) is a family of compact and efficient language models developed by Hugging Face, available in three sizes: $135$M, $360$M, and $1.7$B parameters. We utilise the $360$M parameter variant in our experiments.
B.3 Hyperparameters
In our experiments, we conduct full fine-tuning using a learning rate of $5\!\times\!10^{-5}$, batch size of $8$ and a dropout rate of $0.1$. For the AdamW optimizer, we use $\beta_{1}$, $\beta_{2}$ and $\epsilon$ of $0.9$, $0.999$ and $10^{-8}$ respectively. Each model was trained for $1$ epoch with the hyperparameters fixed for all experiments. We conducted $10$ iterations of recursive training.
Appendix C Informed sampling
C.1 Bias term
To correct for the bias in the classifier’s confidence estimates toward predicting text as machine-generated, we introduce a bias factor $b$ based on the classification threshold:
$b=1+\frac{\text{threshold}}{1-\text{threshold}}.$ (S1)
This formulation is motivated by the odds ratio at the decision boundary, where $\frac{\text{threshold}}{1-\text{threshold}}$ represents the relative likelihood of a sample being classified as machine-generated versus human-generated at the threshold. By adding $1$, we ensure that $b\!\geq 1$, maintaining a minimum correction factor of 1. For a classifier with $\text{threshold}>0.5$, this formulation increases the selection probability of underrepresented samples, counteracting the bias introduced by the classifier’s skewed confidence distribution.
GPT-2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SmolLM2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure S3: GPT-2 (top) and SmolLM2 (bottom) under partially synthetic recursive training with cross-generational data $\left(\alpha\!=\!0.5,\beta\!=\!0.5,\gamma\!=\!0.5\right)$ for 10 generations. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect AI text detector that filters all synthetic samples.
@@ C.1 Bias term
Gen 0 Gen 9 Gen 0 Gen 9 Gen 0 Gen 9 Gen 0 Gen 9 Gen 0 Gen 9 Gen 0 Gen 9
GPT-2 baseline top-$k$5, 1, 0 $29.23$ $31.85$ $38.80$ $37.56$ $84.46$ $82.16$ $39.10$ $39.55$ $93.31$ $91.10$ $51.30$ $46.16$
.5, .5, .5 $29.22$ $31.33$ $38.83$ $37.84$ $85.66$ $84.35$ $38.94$ $40.64$ $94.07$ $92.00$ $51.06$ $45.38$
1, 1, 0 $29.25$ $29.92$ $38.78$ $38.34$ $84.66$ $82.68$ $39.19$ $40.20$ $93.24$ $92.69$ $50.99$ $48.09$
sampling .5, 1, 0 $29.25$ $32.27$ $38.78$ $38.01$ $94.86$ $97.77$ $24.09$ $14.52$ $91.08$ $46.87$ $40.64$ $23.69$
.5, .5, .5 $29.26$ $31.82$ $38.77$ $38.03$ $94.95$ $97.50$ $24.15$ $17.02$ $90.07$ $76.01$ $41.02$ $29.08$
1, 1, 0 $29.25$ $38.79$ $38.79$ $38.63$ $94.88$ $96.86$ $24.04$ $18.12$ $91.18$ $75.64$ $40.71$ $31.49$
ours top-$k$5, 1, 0 $29.25$ $29.53$ $38.78$ $38.45$ $84.43$ $85.89$ $39.43$ $38.85$ $94.59$ $91.52$ $50.79$ $50.33$
.5, .5, .5 $29.25$ $29.47$ $38.77$ $38.49$ $85.18$ $86.40$ $39.28$ $38.90$ $94.75$ $95.67$ $51.31$ $49.99$
1, 1, 0 $29.25$ $28.59$ $38.77$ $38.91$ $84.81$ $85.65$ $38.85$ $38.76$ $94.92$ $93.95$ $51.28$ $51.34$
sampling .5, 1, 0 $29.25$ $29.88$ $38.79$ $38.58$ $94.87$ $96.50$ $24.30$ $19.85$ $92.92$ $81.58$ $40.68$ $35.59$
.5, .5, .5 $29.26$ $29.74$ $38.76$ $38.60$ $94.92$ $96.46$ $24.09$ $20.66$ $91.16$ $88.46$ $40.98$ $36.48$
1, 1, 0 $29.24$ $28.84$ $38.78$ $39.00$ $94.61$ $96.17$ $24.01$ $21.87$ $91.76$ $88.06$ $40.78$ $38.10$
SmolLM2 baseline top-$k$5, 1, 0 $13.36$ $14.06$ $47.89$ $47.57$ $82.50$ $73.99$ $48.99$ $52.77$ $94.58$ $91.10$ $55.65$ $58.09$
.5, .5, .5 $13.36$ $14.00$ $47.89$ $47.58$ $82.46$ $76.24$ $48.66$ $52.44$ $93.68$ $95.13$ $55.24$ $57.88$
1, 1, 0 $13.36$ $13.49$ $47.88$ $47.85$ $82.86$ $77.80$ $48.68$ $51.36$ $94.76$ $89.06$ $55.43$ $57.27$
temperature .5, 1, 0 $13.36$ $14.07$ $47.89$ $47.70$ $83.42$ $66.25$ $45.57$ $52.77$ $95.37$ $81.27$ $52.72$ $57.67$
.5, .5, .5 $13.36$ $13.95$ $47.89$ $47.67$ $83.27$ $71.22$ $45.89$ $52.73$ $88.33$ $84.54$ $53.01$ $56.23$
1, 1, 0 $13.36$ $13.46$ $47.87$ $47.93$ $82.73$ $72.75$ $45.69$ $52.21$ $94.57$ $87.55$ $52.88$ $56.39$
ours top-$k$5, 1, 0 $13.36$ $13.56$ $47.88$ $47.76$ $82.21$ $81.77$ $49.04$ $49.91$ $94.96$ $93.19$ $55.49$ $56.64$
.5, .5, .5 $13.36$ $13.51$ $47.87$ $47.83$ $82.56$ $81.86$ $48.83$ $49.62$ $94.55$ $94.46$ $55.33$ $56.55$
1, 1, 0 $13.36$ $13.15$ $47.88$ $48.08$ $82.35$ $82.52$ $45.00$ $48.91$ $94.00$ $94.43$ $55.35$ $56.14$
temperature .5, 1, 0 $13.36$ $13.64$ $47.88$ $47.79$ $82.83$ $79.35$ $46.08$ $49.15$ $93.13$ $70.63$ $52.87$ $56.08$
.5, .5, .5 $13.36$ $13.62$ $47.88$ $47.80$ $83.31$ $79.62$ $45.64$ $49.34$ $93.07$ $96.14$ $52.79$ $55.68$
1, 1, 0 $13.36$ $13.26$ $47.87$ $48.06$ $83.24$ $80.87$ $45.94$ $48.50$ $93.93$ $78.36$ $52.64$ $54.90$
@@
Table S1: Test performance (perplexity and accuracy) and data quality at generation 00 and generation $9$. Results are shown for top-$k$ decoding and sampling/temperature for different values of $\alpha$, $\beta$, and $\gamma$ ($\uparrow$ / $\downarrow$: higher / lower is better).